OpenStudy (anonymous):

Find the length of the shortest leg of a right triangle which has an hypotenuse of length 15 ft, and one angle of 30 degrees. Find the length of the shortest leg of a right triangle which has an hypotenuse of length 15 ft, and one angle of 30 degrees. @Mathematics

OpenStudy (anonymous):

\[15\times \cos(30) or 15\times \sin(30)\] the smallest one is the shortest leg